# Receding-Horizon Maximum-Likelihood Estimation of Neural-ODE Dynamics and Thresholds from Event Cameras

### 05 Mar 2026 v1

@@ Abstract
Event cameras emit asynchronous brightness-change events where each pixel triggers an event when the last event exceeds a threshold, yielding a history-dependent measurement model. We address online maximum-likelihood identification of continuous-time dynamics from such streams. The latent state follows a Neural ODE and is mapped to predicted log-intensity through a differentiable state-to-image model. We model events with a history-dependent marked point process whose conditional intensity is a smooth surrogate of contrast-threshold triggering, treating the contrast threshold as an unknown parameter. The resulting log-likelihood consists of an event term and a compensator integral. We propose a receding-horizon estimator that performs a few gradient steps per update on a receding horizon window. For streaming evaluation, we store two scalars per pixel (last-event time and estimated log-intensity at that time) and approximate the compensator via Monte Carlo pixel subsampling. Synthetic experiments demonstrate joint recovery of dynamics parameters and the contrast threshold, and characterize accuracy–latency trade-offs with respect to the window length.
@@ I Introduction
Event cameras [5], including the Dynamic Vision Sensor (DVS) [12] and the DAVIS family [1], report asynchronous brightness-change events with microsecond timestamps, high dynamic range, and low latency. Unlike frame-based sensors, they transmit information only when brightness at each pixel changes sufficiently, thereby emphasizing moving edges and reducing motion blur caused by exposure time integration. These properties make event streams well-suited in high-speed and high-dynamic-range settings, enabling optical flow [24, 8, 9]; high-frame-rate video reconstruction and frame interpolation [17, 19, 15, 22]; and ego-motion/SLAM datasets and simulators [14, 16]. Recent work also demonstrates that event cameras can reduce perceptual latency in safety-critical perception pipelines [7].
@@
Many event-based pipelines target a single-shot prediction from a fixed spatiotemporal event representation. In contrast, a broad class of applications requires continuous-time state and dynamics estimates for tracking, prediction, online system identification, and feedback control. In these settings, the information needed for estimation is encoded not only in the event marks (pixel and polarity) but also in fine event timing. This motivates estimators that (i) respect asynchronous timestamps and (ii) provide a likelihood over the raw event stream, so that dynamics parameters can be fit and adapted online in a principled way.
A common description of event generation is contrast thresholding: a pixel fires when the accumulated change in log-intensity since its previous event reaches a threshold. Because the threshold directly affects event timing, it directly impacts dynamics estimation from timestamps. In practice, the effective threshold is often not known precisely and can depend on sensor settings and operating conditions [21, 23]. Treating the threshold as a fixed known constant can therefore introduce bias, especially when the inference relies on fine timing structure.
@@
Beyond threshold uncertainty, event streams pose fundamental modeling challenges. Event times are irregular, observations are discrete marks (pixel location and polarity), and event generation is history-dependent because each pixel resets its reference after it fires. A common workaround is to aggregate events into fixed-rate tensors (for example, voxel grids) and apply standard architectures [24, 17, 25]. Such representations are effective for downstream prediction but discretize time and typically do not define a normalized likelihood over the original timestamps. Complementary model-based approaches incorporate event generation mechanisms for reconstruction and filtering, or optimize surrogate objectives such as contrast maximization and alignment under a motion hypothesis [19, 6, 20, 10]. However, these approaches are usually not formulated as maximum-likelihood estimation of continuous-time dynamics directly from the raw event stream.
@@
A likelihood-based alternative is to model an event stream as a marked temporal point process. This provides a normalized probability model over both event times and marks, and it naturally accommodates irregular timestamps and history dependence. A practical challenge is computational cost. Point-process likelihoods include, in addition to a sum over observed events, a normalization (no-event) term that integrates the predicted event rate over time and sums over the full pixel grid. Evaluating this term can dominate runtime especially for long streams, which is particularly problematic in online estimation settings.
In this paper, we develop an online maximum-likelihood estimator for continuous-time dynamics observed through event-camera streams. Its latent state follows a Neural ODE [2] and is mapped to predicted log-intensity through a differentiable state-to-image model. Events are modeled with a history-dependent marked point process whose conditional intensity is a smooth surrogate of contrast-threshold triggering. We treat the pixel-dependent contrast thresholds as unknown parameters and estimate them jointly with the dynamics. To support streaming operation, we maintain a compact per-pixel memory that summarizes the information needed to evaluate history-dependent residuals. We use a fixed-lag receding-horizon scheme that replays only recent events aiming to keep per-update computation bounded.
@@
Related work: Our work connects three lines of research.
Event-to-tensor pipelines. Many approaches aggregate events into fixed-rate tensors (e.g., voxel grids) and apply standard deep architectures for tasks such as optical flow and video reconstruction [24, 17]. These methods are effective for downstream prediction. However, the aggregation discretizes time and typically does not yield a normalized likelihood over the raw event timestamps. As a result, they are not directly suited to maximum-likelihood identification of continuous-time dynamics (and thresholds) from event streams, especially in an online setting where parameters must be adapted from timing information.
Model-based and surrogate-objective methods. Complementary approaches incorporate event-generation mechanisms for reconstruction and filtering, or optimize surrogate objectives such as contrast maximization and alignment under a motion hypothesis [19, 6, 20, 10]. However, these methods typically optimize task-specific surrogates (e.g., alignment/contrast) or reconstruction losses rather than a point-process likelihood over event times and marks. Moreover, threshold uncertainty is often handled by calibration or fixed settings [21, 23], whereas we treat the thresholds as learnable parameters that are estimated jointly with the dynamics within a single likelihood-based framework.
Temporal point processes and continuous-time neural dynamics. Neural temporal point-process models such as RMTPP [4] and Neural Hawkes [13] provide flexible likelihoods for marked sequences, while continuous-time neural models such as Neural ODEs and latent/jump variants provide convenient representations of dynamics under irregular observations [2, 18, 11]. A direct application of generic point-process training to event cameras is challenging because the mark space is extremely large (pixel locations and polarities) and the observation model is per-pixel history dependent due to the reset at each firing event. Our work exploits this structure explicitly by (i) defining the intensity through a contrast-threshold residual computed from a Neural ODE state via a differentiable state-to-image model, and (ii) making online maximum-likelihood updates practical using a compact per-pixel memory and a fixed-lag receding-horizon scheme with bounded per-update computation.
@@ V-A3 Event synthesis and pixel-dependent threshold
V-B Estimation task
For the dynamics parameters, we aim to estimate
$\vartheta=\{\alpha,\omega\}.$ (34)
@@ V-C1 Learning curves with a fixed horizon ($H=15$)
V-C2 Horizon ablation ($H=1,\dots,19$)
@@
Figure 8: Horizon ablation. Threshold-map error $\mathrm{RMSE}_{C}(H)$ over active pixels (38) as a function of $H$. Note that the y-axis is shown in linear scale for readability.
@@ VI Conclusion and Future Work
We presented a likelihood-based estimator for continuous-time dynamics from event-camera streams by combining a Neural ODE state model with a history-dependent marked point-process observation model. The event generation is modeled via a differentiable surrogate of contrast-threshold triggering, and the (pixel-dependent) contrast threshold is treated as an unknown parameter to be estimated jointly with the dynamics. To enable streaming maximum-likelihood training, we approximate the compensator term using Monte Carlo pixel subsampling, and summarize the required per-pixel history with a compact boundary memory (two scalars per pixel). A fixed-lag receding-horizon update then bounds per-update computation by restricting optimization to a sliding time window.

### 03 Apr 2026 v2

@@ Abstract
Event cameras produce asynchronous brightness-change events when the accumulated log-intensity change at a pixel since its previous event reaches a threshold, yielding a history-dependent observation process. We study online maximum-likelihood identification of continuous-time dynamics directly from such raw event streams. The latent state evolves according to a Neural ODE and is mapped to predicted log-intensity through a differentiable state-to-image model. We model the event stream as a history-dependent marked temporal point process whose conditional intensity is a smooth probabilistic surrogate of contrast-threshold triggering, and we treat the pixel-dependent contrast thresholds as unknown parameters to be estimated jointly with the dynamics. This yields a normalized likelihood over event times and marks consisting of an event term and a compensator integral. To enable streaming inference, we develop a fixed-lag receding-horizon estimator with compact per-pixel boundary memory and Monte Carlo pixel subsampling for the compensator. Synthetic experiments demonstrate joint recovery of the dynamics parameters and the contrast threshold map, and quantify the resulting accuracy–latency trade-off as the horizon length varies.
@@ I Introduction
Event cameras [12], including the Dynamic Vision Sensor (DVS) [20] and the DAVIS family [2], report asynchronous brightness-change events with microsecond timestamps, high dynamic range, and low latency. Unlike frame-based sensors, they transmit information only when brightness at each pixel changes sufficiently, thereby emphasizing moving edges and reducing motion blur caused by exposure time integration. These properties make event streams well-suited to high-speed and high-dynamic-range settings, enabling optical flow [44, 14, 15]; high-frame-rate video reconstruction and frame interpolation [29, 32, 25, 40]; and ego-motion/SLAM datasets and simulators [24, 28]. Recent work further highlights event-driven sensing and event-camera processing for efficient perception, video reconstruction, and stereo depth estimation [21, 17, 6].
@@
Many event-based pipelines target a single-shot prediction from a fixed spatiotemporal event representation. In contrast, a broad class of applications requires continuous-time state and dynamics estimates for tracking, prediction, online system identification, and feedback control. In these settings, the information needed for estimation is encoded not only in the event marks (pixel and polarity) but also in fine event timing. This motivates estimators that (i) respect asynchronous timestamps and (ii) provide a likelihood over the raw event stream, so that dynamics parameters can be fit and adapted online in a principled way. Asynchronous thresholded observations have been studied in time encoding, level-crossing sampling, unlimited sampling, and spike-train reconstruction, where information is represented by event times rather than uniform samples [19, 34, 3, 11, 4]. In parallel, statistical modeling and inference for event times has continued to develop in temporal point-process and event-triggered estimation settings [37, 36, 10, 43, 22, 30, 8]. Event-camera streams combine both aspects: each pixel generates asynchronous thresholded observations, while the full sensor output is naturally a marked event sequence whose information is carried jointly by event times and marks.
@@
A common description of event generation is contrast thresholding: a pixel fires when the accumulated change in log-intensity since its previous event reaches a threshold. Because the threshold directly affects event timing, it directly impacts dynamics estimation from timestamps. In practice, the effective threshold is often not known precisely and can depend on sensor settings and operating conditions [39, 41]. Treating the threshold as a fixed known constant can therefore introduce bias, especially when the inference relies on fine timing structure. Beyond threshold uncertainty, event streams pose fundamental modeling challenges. Event times are irregular, observations are discrete marks (pixel location and polarity), and event generation is history-dependent because each pixel resets its reference after it fires. A common workaround is to aggregate events into fixed-rate tensors (for example, voxel grids) and apply standard architectures [44, 29, 17, 6, 46]. Such representations are effective for downstream prediction but discretize time and typically do not define a normalized likelihood over the original timestamps. Complementary model-based approaches incorporate event-generation mechanisms for reconstruction and filtering, or optimize surrogate objectives such as contrast maximization and alignment under a motion hypothesis [32, 13, 38, 16]. However, these approaches are usually not formulated as maximum-likelihood estimation of continuous-time dynamics directly from the raw event stream.
@@
A likelihood-based alternative is to model an event stream as a marked temporal point process. This provides a normalized probability model over both event times and marks, and it naturally accommodates irregular timestamps and history dependence. A practical challenge is computational cost. Point-process likelihoods include, in addition to a sum over observed events, a normalization (no-event) term that integrates the predicted event rate over time and sums over the full pixel grid. Evaluating this term can dominate runtime especially for long streams, which is particularly problematic in online estimation settings. Sliding-window and moving/receding-horizon estimators, as well as recent event-triggered online estimation methods, provide a natural route to bounded-latency updates in dynamical systems [26, 27, 1, 35, 45, 43, 22]. For event cameras, however, the observation model is both high-dimensional and per-pixel history dependent, which makes a direct application of existing methods nontrivial.
In this paper, we develop an online maximum-likelihood estimator for continuous-time dynamics observed through event-camera streams. Its latent state follows a Neural ODE [5] and is mapped to predicted log-intensity through a differentiable state-to-image model. Events are modeled with a history-dependent marked point process whose conditional intensity is a smooth surrogate of contrast-threshold triggering. We treat the pixel-dependent contrast thresholds as unknown parameters and estimate them jointly with the dynamics. To support streaming operation, we maintain a compact per-pixel memory that summarizes the information needed to evaluate history-dependent residuals. We use a receding-horizon scheme that replays only recent events, aiming to keep per-update computation bounded.
@@
Related work: Our work connects four lines of research.
Event-driven sensing and asynchronous threshold acquisition. Recent work studies event-driven sensing and event-camera processing for efficient perception, video reconstruction, and stereo depth estimation [21, 17, 6]. Closely related time-encoding and level-crossing/unlimited-sampling work studies how continuous-time signals can be represented and recovered from threshold-crossing times [19, 34, 3, 11, 4]. Event cameras can be viewed as a spatially distributed, polarity-marked extension of this principle. Unlike classical time-encoding setups, however, the quantity of interest here is not the per-pixel brightness trace itself but latent continuous-time dynamics observed through a state-to-image map, and the thresholds are not assumed known.
Event-camera pipelines and model-based objectives. Many approaches aggregate events into fixed-rate tensors (e.g., voxel grids) and apply standard deep architectures for tasks such as optical flow and video reconstruction [44, 29, 17, 6]. These methods are effective for downstream prediction. However, the aggregation discretizes time and typically does not yield a normalized likelihood over the raw event timestamps. Complementary approaches incorporate event-generation mechanisms for reconstruction and filtering, or optimize surrogate objectives such as contrast maximization and alignment under a motion hypothesis [32, 13, 38, 16]. In particular, the spatio-temporal Poisson point-process model in [16] is closer in spirit because it defines a likelihood over event data, but it is aimed at event alignment under a motion hypothesis rather than joint estimation of continuous-time latent dynamics and pixel-dependent thresholds from a per-pixel history-dependent observation model.
Statistical inference with point-process or event-triggered observations. Classical work on counting/point-process observations studies filtering and state-space inference for latent dynamical systems observed through event times [37, 33, 36, 10, 42]. Recent work further studies point-process learning, classification, and event-triggered online estimation [30, 8, 43, 22]. In a related vein, moving-horizon estimation with binary or threshold sensors has been developed for lower-dimensional dynamical systems [26, 27, 1, 35, 45]. Our setting differs in that the mark space is extremely large (pixel location and polarity), the observation model is per-pixel history dependent due to the reset at each firing event, and the thresholds are learned jointly with the dynamics.
Continuous-time neural dynamics. Neural temporal point-process models such as RMTPP [9] and Neural Hawkes [23] provide flexible likelihoods for marked sequences, while continuous-time neural models such as Neural ODEs and latent/jump variants provide convenient representations of dynamics under irregular observations [5, 31, 18]. Our work exploits the event-camera structure explicitly by (i) defining the intensity through a contrast-threshold residual computed from a Neural ODE state via a differentiable state-to-image model, and (ii) making online maximum-likelihood updates practical using a compact per-pixel memory and a fixed-lag receding-horizon scheme with bounded per-update computation.
@@ V-A3 Event synthesis and pixel-dependent threshold
V-B Estimation task
For the dynamics parameters, we aim to estimate
$\vartheta=\{\alpha,\omega\}.$ (34)
Figure 6: Snapshots of pixel-dependent contrast-threshold estimation over online episodes ($H=15$). Left: ground-truth threshold field $C(u)$ used for event synthesis. Right: estimated maps $\hat{C}(u)$ after episodes $e\in\{1,17,32\}$.
@@ V-C1 Learning curves with a fixed horizon ($H=15$)
V-C2 Horizon ablation ($H=1,\dots,34$)
@@
Figure 8: Horizon ablation. Threshold-map error $\mathrm{RMSE}_{C}(H)$ over active pixels (38) as a function of $H$. Note that the y-axis is shown in linear scale.
@@ VI Conclusion and Future Work
We presented an online maximum-likelihood framework for identifying continuous-time dynamics from raw event-camera streams. By combining a Neural ODE state model with a history-dependent marked point-process observation model, the proposed formulation captures the asynchronous thresholded nature of event generation while preserving a normalized likelihood over event times and marks. A differentiable surrogate of contrast-threshold triggering enables joint estimation of the dynamics and pixel-dependent thresholds, and the fixed-lag receding-horizon update with compact boundary memory and Monte Carlo compensator approximation keeps the per-update computation bounded. On synthetic data, the method recovers both the dynamics parameters and the threshold map and reveals the accuracy–latency trade-off induced by the horizon length. These results suggest that likelihood-based continuous-time identification from raw event timing is a viable alternative to event aggregation when online adaptation is required.
Future work includes validation on real event-camera data, incorporation of richer sensor nonidealities such as polarity-dependent thresholds and temporal drift, and improvement of the compensator approximation via adaptive sampling. Extending the framework to joint online state-and-parameter estimation in more complex scenes and studying uncertainty and identifiability are also important directions for future work.